First and Ten Corporation’s stock returns have a covariance with the market portfolio of .0516. The standard deviation of the returns on the market portfolio is 21 percent and the expected market risk premium is 6.3 percent. The company has bonds outstanding with a total market value of $57 million and a yield to maturity of 6.6 percent. The company also has 6.2 million shares of common stock outstanding, each selling for $31. The company’s CEO considers the firm’s current debt-equity ratio optimal. The corporate tax rate is 25 percent and Treasury bills currently yield 4.6 percent. The company is considering the purchase of additional equipment that would cost $59 million. The expected unlevered cash flows from the equipment are $19.4 million per year for 5 years. Purchasing the equipment will not change the risk level of the firm. Calculate the NPV of the project. (Do not round intermediate calculations and enter your answer in dollars, not millions of dollars, rounded to 2 decimal places, e.g., 1,234,567.89)
Answer #1
1] Beta of the stock = Covariance of the stock returns with the market portfolio/Variance of the market portfolio =
=0.0516/0.21^2 = 1.17
2] Cost of equity per CAPM = 4.6%+1.17*6.3% = 11.97%
3] After tax cost of debt = 6.6%*(1-25%) = 4.95%
4] Calculation of WACC:
Source of capital Market value in $ million Weight Component cost WACC
Debt [MV given] $                    57.00 22.87% 4.95% 1.13%
Common stock [MV = 6.2*31] $                  192.20 77.13% 11.97% 9.23%
Total $                  249.20 10.37%
WACC = 10.37%
5] NPV = -59000000+19400000*(1.1037^5-1)/(0.1037*1.1037^5) = $ 1,38,51,375.60